We call a positive integer perfect if it equals the sum of its positive divisors


other than itself.


(a) Prove that 6 and 28 are perfect numbers


(b) Prove that if 2p − 1 is prime, then 2p−1



(2p − 1) is a perfect number


1
Expert's answer
2021-11-03T10:03:49-0400

Question is not written in proper way.

Actual question is:

We call a positive integer perfect if it equals the sum of its positive divisors other than itself.

a) Prove that 6 and 28 are perfect.

b) Prove that "2^{p\u22121}(2^p \u2212 1)" is a perfect number when "2^p \u22121" is prime.

Solution:

(a) Let us determine all divisors of each number.

Divisors of "6=1,2,3,6"

Divisors of "28=1,2,4,7,14,28"

Let us determine the sum of all positive divisors other than the integer itself:

"\begin{gathered}

6=1+2+3 \\

28=1+2+4+7+14

\end{gathered}"

We then note that 6 and 28 are both perfect.

(b) Given: "2^{p}-1" is prime

To prove: "2^{p-1}\left(2^{p}-1\right)" is a perfect number

Poof:

Determine all divisors of "2^{p-1}\left(2^{p}-1\right)" (using that "\left(2^{p}-1\right)" is prime and thus cannot be factorized):

"1,2,2^{2}, \ldots, 2^{p-1},\left(2^{p}-1\right), 2\left(2^{p}-1\right), 2^{2}\left(2^{p}-1\right), \ldots .2^{p-2}\left(2^{p}-1\right), 2^{p-1}\left(2^{p}-1\right)"

Note: The last divisor is the number "2^{p-1}\left(2^{p}-1\right)" itself.

Let us determine the sum of all positive divisors other than the integer itself:

"\begin{aligned}
&1+2+2^{2}+\ldots+2^{p-1}+\left(2^{p}-1\right)+2\left(2^{p}-1\right)+2^{2}\left(2^{p}-1\right)+\ldots+2^{p-2}\left(2^{p}-1\right) \\
&=1\left(1+2^{p}-1\right)+2\left(1+2^{p}-1\right)+2^{2}\left(1+2^{p}-1\right)+\ldots+2^{p-2}\left(1+2^{p}-1\right)+2^{p-1} \\
&=1\left(2^{p}\right)+2\left(2^{p}\right)+2^{2}\left(2^{p}\right)+\ldots+2^{p-2}\left(2^{p}\right)+2^{p-1} \\
&=2^{p}\left(1+2+2^{2}+\ldots+2^{p-2}\right)+2^{p-1}
\end{aligned}"

"\begin{aligned}
&=2^{p}\left(\sum_{i=0}^{p-2} 2^{i}\right)+2^{p-1} \\
&=2^{p}\left(\frac{2^{p-1}-1}{2-1}\right)+2^{p-1} \\
&=2^{p}\left(2^{p-1}-1\right)+2^{p-1} \\
&=2^{p} \cdot 2^{p-1}-2^{p}+2^{p-1} \\
&=2^{p} \cdot 2^{p-1}-2 \cdot 2^{p-1}+1 \cdot 2^{p-1} \\
&=\left(2^{p}-2+1\right) 2^{p-1} \\
&=\left(2^{p}-1\right) 2^{p-1} \\
&=2^{p-1}\left(2^{p}-1\right)
\end{aligned}"

By the definition of a perfect number, we have then shown that "2^{p-1}\left(2^{p}-1\right)" is a perfect number.

Hence Proved